OpenStudy (yash2651995):

draw the graph of: sinx / (1+ (tanx^2))^(1/2) + cosx / (1+ (cotx^2))^(1/2)

OpenStudy (anonymous):

A plot of\[\frac{\sin (x)}{\sqrt{\tan ^2(x)+1}}+\frac{\cos (x)}{\sqrt{\cot ^2(x)+1}} \]from x=0 through x=2Pi is attached.

OpenStudy (yash2651995):

thanks a lot :)